OpenStudy (anonymous):

Given b=4 and h=1, what is the equation of the graph if the parent function is y=√x

OpenStudy (anonymous):

what are b and h lol

OpenStudy (anonymous):

b=4 and h=1 lol

OpenStudy (anonymous):

what b and h stands for

OpenStudy (anonymous):

oh idk.. Just gives me options

OpenStudy (anonymous):

uh the things not working lol Hold on i'm trying to put answers uop

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

\[\sqrt{4x-4}\] \[\sqrt{4x+4}\] \[-\sqrt{4x-4}\] \[-\sqrt{4x+4}\]

OpenStudy (anonymous):

i think its a , 90% sure bout it

OpenStudy (anonymous):

lol thanks haha

OpenStudy (anonymous):

np ;)

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

;)

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

haha